OpenStudy (anonymous):

If it takes 37 minutes 58 seconds to complete a task, how long would it take to repeat the same task 7 times (answer in hours, minutes and seconds)?

OpenStudy (anonymous):

I got the hours but I am finding it hard getting the seconds

OpenStudy (anonymous):

if you have the hours, just multiply it by 3600 to get the seconds!

OpenStudy (anonymous):

No thats that the answer, its meant to be in hours, minutes and seconds. I have found the hours part just cant change it to the rest

OpenStudy (anonymous):

well.. here is the answer.. the time taken to complete one task, in seconds is 37*60+58=2278sec the time to do 7 tasks 7*2278=15946sec.

OpenStudy (anonymous):

I got all of that then divided bY 3600

OpenStudy (anonymous):

the time in minutes=15946/60=265.77 min the time in hours=4.43 hours

OpenStudy (anonymous):

so what do you exactly want?

OpenStudy (anonymous):

this is the time it takes to do 7 tasks, in hours, minutes and seconds

OpenStudy (anonymous):

No u dont get it. these are the answer choice a 4 hours 26 mins 0 secs b 3 hours 45 mins 46 secs c 4 hours 25 mins 46 secs d 4 hours 66 mins 15 secs

OpenStudy (anonymous):

I see

OpenStudy (amistre64):

(37 min*7) + (58sec*7) 259min + 406 sec ..... maybe :)

OpenStudy (anonymous):

O.O that looks nice actually

OpenStudy (amistre64):

259/60 = 4hrs and 16 minutes

OpenStudy (anonymous):

ok.. 4.43 hours=4 hours +0.43*60 min=4 hours 25.8 min=4 hours 25min 48sec

OpenStudy (amistre64):

406/60 = 6minutes and 46 seconds i think

OpenStudy (amistre64):

4 hours 22 minues and 46 seconds if I did it right... which is a BIG if today :)

OpenStudy (anonymous):

the difference is due to the approximation I made in an earlier step :)

OpenStudy (anonymous):

the answer is c

OpenStudy (anonymous):

anistre64 your answer is not in the options a 4 hours 26 mins 0 secs b 3 hours 45 mins 46 secs c 4 hours 25 mins 46 secs d 4 hours 66 mins 15 secs Ok got it now.. Thank you

OpenStudy (amistre64):

its close... accounting for errors in my stupidity :)

OpenStudy (anonymous):

it's exactly c.. I double checked it

OpenStudy (anonymous):

yep XD. Thanks again

OpenStudy (amistre64):

259/60 was off a bit...4 hours 19 minutes

OpenStudy (amistre64):

4.19.0 + 0.6.46 4.25.46 is right...

OpenStudy (anonymous):

nice, you way seems shorter and easier as well. thanks

OpenStudy (anonymous):

Kadious, do you know how to do it anyway? for next times? :P

OpenStudy (amistre64):

i think its the same, just constructed differently,,,, :)

OpenStudy (anonymous):

yep, I got up the 4.43 hours bit but didnt know how to get the mins and seconds

OpenStudy (anonymous):

But now I got the idea. Basically to multiply the 0.43 by 60 then if there are any remainder, I mulitply by 60 again

OpenStudy (amistre64):

instead of going on into the decimals, keep the remainder of the division to "see" whats left :)

OpenStudy (anonymous):

Hmm .. sounds and seems harder.

OpenStudy (amistre64):

another way is to "estimate" the time to 38 minutes; then subtract from that 7*2secs ...... 38 * 7 = 266 minutes 266/60 = 4hrs with a remainder of 26minutes -7*2 secs = -14 sec 4 hrs 25 min and 46 sec

OpenStudy (anonymous):

oh thats nice. so basically the the 2 is the 60-58 and 7 is the "7 times" thingy. So its basically the same as the other one you did

OpenStudy (amistre64):

yep..

OpenStudy (anonymous):

nice , Thanks XD